Two exterior angles of a polygon are 63° each.The remaining exterior angles are each 26°.How many sides has the polygon?
Ive thought about it, please help.
Thanks

Answer by Alan3354(69443)

Two exterior angles of a polygon are 63° each.The remaining exterior angles are each 26°.How many sides has the polygon?
-------------
The sum of the exterior angles for all polygons is 360 degs
360 = 63*2 + x*26
26x = 360 - 126 = 234
x = 9
--> 11 sides